# Generating and Understanding Tests via Path-Aware Symbolic Execution with LLMs

### v1

@@ Abstract
Symbolic execution is a widely used technique for test generation, offering systematic exploration of program paths through constraint solving. However, it is fundamentally constrained by the capability to model the target code including library functions in terms of symbolic constraint and the capability of underlying constraint solvers. As a result, many paths involving complex features remain unanalyzed or insufficiently modeled.
Recent advances in large language models (LLMs) have shown promise in generating diverse and valid test inputs. Yet, LLMs lack mechanisms for systematically enumerating program paths and often fail to cover subtle corner cases. We observe that directly prompting an LLM with the full program leads to missed coverage of interesting paths.
@@
In this paper, we present PALM, a test generation system that combines symbolic path enumeration with LLM-assisted test generation. PALM statically enumerates possible paths through AST-level analysis and transforms each into an executable variant with embedded assertions that specify the target path. This avoids the need to translate path constraints into SMT formulae, by instead constructing program variants that LLM can interpret.
Importantly, PALM is the first to provide an interactive frontend that visualizes path coverage alongside generated tests, assembling tests based on the specific paths they exercise. A user study with 12 participants demonstrates that PALM’s frontend helps users better understand path coverage and identify which paths are actually exercised by PALM-generated tests, through verification and visualization of their path profiles.
I Introduction
@@
Symbolic execution is a widely adopted software testing and verification technique [1, 2]. It systematically explores different program paths and collects the branch predicates along each path as logical constraints, a.k.a. path constraints. These constraints are typically encoded as SMT formulas and passed to constraint solvers for concrete test input generation.
@@
However, symbolic execution is fundamentally limited by the symbolic encoding required for individual functions and subsequent constraint solving. As shown in Fig. 1, the method parseFilePath processes a list of arguments and attempts to extract the file path following the "-f" flag. A key branch condition in this method involves the string comparison args[i].equalsIgnoreCase("-f"). Constraint solvers such as Z3[3] or CVC5 [4] lack built-in support for such operations, unless symbolic execution tool developers manually supply symbolic models for the corresponding library function, such as equalsIgnoreCase.
@@
Recent studies [5, 6] have demonstrated that LLM-assisted testing is very effective in automated test generation. LLMs are capable of understanding program semantics and producing test inputs.
@@
Second, existing LLM-based test generation tools primarily focus on maximizing overall code coverage through automated test generation, but offer little support for understanding which specific execution paths are exercised by each generated test. To the best of our knowledge, none of them provide visibility into which branches (or paths) remain uncovered or visualize whether generated tests are sufficient in terms of path coverage, limiting the user’s ability to identify missing coverage or guide the generation process toward specific behaviors of interest.
@@ 4 String path = null;
17}
18// GPT-4o generated tests
19parseFilePath(new String[]{});
20parseFilePath(new String[]{"-v"});
21parseFilePath(new String[]{"-f", "input.txt"});
22parseFilePath(new String[]{"-v", "-f", "data.csv"});
23parseFilePath(new String[]{"-f", "log.txt", "-v"});
24parseFilePath(new String[]{"-f"});
Figure 1: This code snippet parses the argument following ”-f” as a file path. While GPT-4o generates tests covering typical cases, it misses the edge case {"-f","-v"}, where "-v" is mistakenly interpreted as the file path due to the lack of validation. This path is also challenging for symbolic execution-based testing as SMT solvers like Z3 and CVC5 do not support string operations such as equalsIgnoreCase.
In this paper, we present PALM, a test generation and comprehension support that combines the path enumeration capability of symbolic execution with the test generation strength of LLMs. PALM also features an interactive frontend that visualizes the symbolic execution tree, highlights path coverage, and allows users to inspect or refine test generation for specific paths.
PALM’s test generation consists of two phases:
@@
PALM performs AST-level analysis on the subject program to systematically enumerate program paths. For each path, it automatically constructs a corresponding program variant using program transformation, which is a modified version of the original program. The variant is augmented with assertTrue and assertFalse statements that encode the branch decisions along the path. This variant serves as a path-specific prompt to guide LLM which test input it must generate. In other words, this variant directly serves as the hint for path-specific test generation.
@@
In our experiments, we evaluated Palm on 124 Java programs from HumanEval-Java [7], which include complex control flows such as nested loops and which use external library functions. By systematically enumerating and extracting program paths, Palm achieves 35.0% and 24.2% higher path coverage than GPT-4o-mini and GPT-o3-mini, respectively, when using the same LLM backend. Its iterative test validation and refinement further improves path coverage by 14.2% over the non-iterative setting with GPT-4o-mini. In contrast, Symbolic PathFinder (a representative traditional symbolic executor) fails to derive a single accurate path constraint in 34.3% of the programs due to insufficient symbolic modeling of external API calls. We also conducted a within-subject user study with 12 participants to understand how much PALM can assist users in generating and comprehending path-aware test inputs. Participants reported higher confidence in their ability to generate sufficient tests, disambiguate redundant or missing tests, and check whether generated tests indeed exercises specific paths, compared to using a pure LLM as is.
@@
  • We conducted a user study with 12 participants, which demonstrates that PALM’s interactive frontend helps users better understand path coverage and identify individual path-exercising tests.
Figure 2: PALM’s user interface. 1 Users provide the subject program under test along with symbolic execution configurations, such as a loop bound and symbolic methods. 4 The symbolic execution tree visualizes all enumerated program paths. Each node represents a program construct: yellow diamonds for branch conditions, blue rectangles for statements, and circles for terminal paths. Green/red/gray circles indicate whether a path is covered, uncovered, or unreachable, respectively. 5 Users can click on any circular node to select a program path, which is then highlighted in orange. 6 Displays the path-specific program variant of the selected path, an equivalent java program representing the execution trace along this path, with inserted assertTrue/assertFalse statements encoding each branching decision. This is passed as an instruction to the LLM for test generation. 7 Shows PALM’s iterative test generation attempts for the selected path. The path exercisability of each generated test is validated with runtime; results are marked green, if successful or red otherwise. 8 Shows the exact prompt given to the LLM for the selected test case. This enables users to analyze and refine prompt quality. 9 Users can compose their own test cases. PALM supports verifying if a given test exercises the selected path or identifying which path it covers.
II Tool Overview
@@
PALM is a test generation system for Java programs that combines symbolic execution with LLMs. For each enumerated path, PALM uses an LLM to generate a path-exercising test input. Its interface centers around a symbolic execution tree, a tree-based representation where each node corresponds to a program statement or a branch condition, and each path from the root to a leaf represents a possible execution path. The symbolic tree, shown in the dashboard, visualizes explored paths and highlights covered ones in green. Users can visualize the execution trace of a given test and align test inputs with path-level program execution behavior. In the following three subsections, we explain PALM’s key features using the example shown in Fig. 2.
II-A Test generation
@@
At its core, PALM systematically enumerates program paths and invokes an LLM to generate test inputs for each enumerated path. Users begin by providing the subject program, such as a program named tutorial in the example figure ( 1), along with symbolic execution configurations. In this example, the method tutorial is marked as the entry point and symbolic, meaning its internal paths will be enumerated. Although the tutorial program contains no loops, users can still configure a loop bound, such as 2, to control the loop unrolling depth in programs. After clicking the Test Generation Button ( 3), PALM begins enumerating paths, constructing program variants, and requesting LLMs for path-specific test generation.
II-B Path understanding with symbolic tree and path variant
@@
An interactive symbolic execution tree serves as a central component in PALM. This tree visualizes the explored execution paths of the program, distinguishing between branch conditions (yellow diamond nodes), regular statements (blue rectangular nodes), and path coverage status—green for covered, red for uncovered, and gray for unreachable paths. For example, the highlighted orange path ( 5) includes two conditional branches, x>0 and y+z>0, both of which must evaluate to true. The non-conditional assignment z=-z-5 appears between them and affects the outcome of the second condition, demonstrating a data dependency along the path.
@@
The variant not only helps users interpret the logic of an execution path, but also serves as a prompt that guides the LLM to generate test inputs satisfying the conditions along that path. Unlike traditional symbolic execution tools that translate path constraints into SMT formulae, PALM leverages LLMs to handle complex language features—such as string operations and library calls—that are often inexpressible in SMT logic. This approach eliminates the need for manually modeling library functions or treating them as uninterpreted, a key advantage over conventional symbolic executors.
II-C Validating and refining generated tests
@@
Users can inspect the divergence and hypothesize the cause. For example, the LLM might ignore the update z=-z-5 before evaluating the y+z>0 branch. To guide the LLM, users can manually edit the prompt in the Prompt Box ( 8) and inject guidance such as: ”The y+z$>$0 branch should consider that z is updated as -z-5.” Users can then update the test to tutorial(1,6,0) and verify whether it covers the intended path ( A), or identify the path actually taken by a test ( B).
@@
III Methodology
Figure 3: PALM operates in two phases: (1) it enumerates and transforms program paths into concrete path-specific program variants through loop unrolling, function inlining, variable renaming, and constant propagation; (2) it traverses the symbolic execution tree, invokes an LLM to generate concrete test inputs and iteratively validates them through runtime execution whether the tests exercise the intended path, regenerating with error feedback if necessary.
@@ III Methodology
PALM takes a Java program as input and systematically enumerates its execution paths using AST-level analysis. For each enumerated path, PALM generates a corresponding path-specific program variant, which is an executable program that represents the statements and branch decisions along the path. These variants encode each path’s semantics in a form of normal Java statements that is interpretable by LLMs, without translating to SMT constraints.
To facilitate precise control over test generation, PALM organizes and visualizes paths into a symbolic execution tree.
@@
The following subsections detail PALM’s two phases.
III-A Path Extraction
@@
4 asssertTrue(i < len); 2 // [1]
5 asssertTrue(text.charAt(i) != // [2]
@@
4 asssertTrue(i < len); 2 // [1]
5 asssertFalse(text.charAt(i) != // [2]
@@
8 asssertTrue(i < len); 5 // [3]
9 asssertTrue(text.charAt(i) != // [4]
@@
Figure 4: This figure illustrates a palindrome-checking program with a for loop and two branching conditions, highlighted in green and yellow, respectively. Subfigure (b) presents the symbolic execution tree, where each branch corresponds to a branching predicate in the source program. By unrolling the loop and resolving branches, we derive concrete path conditions represented as assertion-based encodings in (c) and (d). Each original branch is translated into an assertTrue or assertFalse over the corresponding predicate. For example, 3 means that the conditional (yellow) evaluates to true in the first loop iteration.
@@
Algorithm 1 formalizes our path enumeration process. It recursively traverses the AST of the subject program and compositionally builds the set of enumerated paths. Lines 2–5 handle If-Then-Else structures by separating the two branches and inserting assertions, assertTrue or assertFalse according to the branch decision. For instance, 4 and 6 correspond to two branches of the inner conditional check text.charAt(i) != text.charAt(len-i-1) in Fig. 4.
Lines 6–8 handle While loops by unrolling the loop body up to a specified loop-unrolling bound $K$. Other loop constructs such as Do-While and For loops are handled analogously. Line 10 processes compound BlockStatement structures consisting of sequential statements by enumerating the combinations of subpaths induced by each statement.
@@
1:Program $program$, loop bound $K$ (default = 2)
2:$Paths_{m}$: A set of paths for the symbolic function $m$
3:
4:function ExtractPath($node$) $\triangleright$ $node$ is an AST node
5:    if $node$ is IfStatement then
6:       $P_{\text{then}}\leftarrow$ prepend assertTrue($node.branch$) to each path in ExtractPath($node.thenBlock$)
7:       $P_{\text{else}}\leftarrow$ prepend assertFalse($node.branch$) to each path in ExtractPath($node.elseBlock$)
8:       return $P_{\text{then}}\cup P_{\text{else}}$
9:    else if $node$ is WhileStatement then
10:       Unroll the loop for 0 to $K$ iterations
11:       Apply ExtractPath to each unrolled version
12:    else if $node$ is BlockStatement then
13:       return combinations of ExtractPath($s$), for all $s\in node$
14:    end if
15:    … $\triangleright$ Handle other code structures
16:end function
A common practice in symbolic execution is to allow annotations specifying which functions should undergo path enumeration—referred to as symbolic functions—as well as which function serves as the entry point. This allows users to focus on path exploration on functions of interest. PALM supports this flexibility as well.
Since path extraction flattens control structures and isolates a single execution path per symbolic function, key semantic information may be lost if not preserved explicitly. To address this, we apply function inlining and variable renaming to preserve the semantics of nested calls and variable scopes. Additionally, we perform constant propagation to simplify each path-specific program variant, reducing syntactic complexity and easing LLM reasoning.
@@
1:$P_{m}$: Set of symbolic paths for each symbolic function $m$
2:$R$: Recursion bound (default = 2)
3:$T_{m}$: Complete set of inlined paths for function $m$
4:$T\leftarrow$ paths without symbolic calls
5:$N\leftarrow$ paths with symbolic calls
6:for recursion level $r=1$ to $R$ do
7:    $N^{\prime}\leftarrow\emptyset$
8:    for each non-terminating path $n\in N$ do
9:       Identify call-sites $c_{1},\dots,c_{k}$ in $n$ calling functions $m_{1},\dots,m_{k}$
10:       for all path tuples $(p_{1},\dots,p_{k})$ where $p_{i}\in T_{m_{i}}$ do
11:          $n^{\prime}\leftarrow$ replace each $c_{i}$ in $n$ with $p_{i}$
12:          $N^{\prime}\leftarrow N^{\prime}\cup\{n^{\prime}\}$
13:       end for
14:    end for
15:    $T\leftarrow T\cup N^{\prime}$triangleright Prepare for next inlining level
16:end for
@@
In the original subject program, variable scopes are managed implicitly by the nesting structure. However, after path extraction with control flow expansion—such as loop unrolling or function inlining—these scopes are flattened, and variable reuse can lead to name conflicts or ambiguity. To preserve semantic clarity and avoid confusion during test generation, PALM renames local variables by appending a numeric suffix.
@@
PALM applies intra-procedural constant propagation and constant folding, similar to standard compiler optimizations. For example, in Fig. 4, 2 contains assertTrue(i<len); where i is initialized to 0; this can be folded into assertTrue(0<len). Constant folding also enables pruning of trivially infeasible paths, such as assertTrue(x < y) where x=1 and y=0. These simplifications reduce the syntactic complexity of path variants, easing both the LLM’s symbolic reasoning and human understanding of path semantics.
III-B Path Aware Test Generation with Validation

### v2

# PALM: Path-aware LLM-based Test Generation with Comprehension

@@ Abstract.
Symbolic execution is a widely used technique for test generation, offering systematic exploration of program paths through constraint solving. However, it is fundamentally constrained by the capability to model the target code including library functions in terms of symbolic constraint and the capability of underlying constraint solvers. As a result, many paths involving complex features remain unanalyzed or insufficiently modeled. Recent advances in large language models (LLMs) have shown promise in generating diverse and valid test inputs. Yet, LLMs lack mechanisms for systematically enumerating program paths and often fail to cover subtle corner cases. We observe that directly prompting an LLM with the full program leads to missed coverage of interesting paths.
@@
In this paper, we present PALM, a test generation system that combines symbolic path enumeration with LLM-assisted test generation. PALM statically enumerates possible paths through AST-level analysis and transforms each into an executable variant with embedded assertions that specify the target path. This avoids the need to translate path constraints into SMT formulas, by instead constructing program variants that LLM can interpret. Importantly, PALM provides an interactive frontend that visualizes path coverage alongside generated tests, assembling tests based on the specific paths they exercise. A user study with 12 participants demonstrates that PALM’s frontend helps users better understand path coverage and identify which paths are actually exercised by PALM-generated tests, through verification and visualization of their path profiles.
LLM-assisted testing, symbolic execution
journalyear: 2026copyright: ccconference: 34th IEEE/ACM International Conference on Program Comprehension; April 12–13, 2026; Rio de Janeiro, Brazilbooktitle: 34th IEEE/ACM International Conference on Program Comprehension (ICPC ’26), April 12–13, 2026, Rio de Janeiro, Brazildoi: 10.1145/3794763.3794797isbn: 979-8-4007-2482-4/2026/04ccs: Software and its engineering Software testing and debugging
1. Introduction
@@
Symbolic execution is a widely adopted software testing and verification technique (King, 1975; Boyer et al., 1975; King, 1976; Howden, 1977). It systematically explores different program paths and collects the branch predicates along each path as logical constraints, a.k.a. path constraints. These constraints are typically encoded as Satisfiability Modulo Theories (SMT) formulas and passed to constraint solvers for concrete test input generation.
@@
However, symbolic execution is fundamentally limited by the symbolic encoding required for individual functions and subsequent constraint solving. As shown in Fig. 1, the method parseFilePath processes a list of arguments and attempts to extract the file path following the "-f" flag. A key branch condition in this method involves the string comparison args[i].equalsIgnoreCase("-f"). Constraint solvers such as Z3 (De Moura and Bjørner, 2008) or CVC5 (Barbosa et al., 2022) lack built-in support for such operations, unless symbolic execution tool developers manually supply symbolic models for the corresponding library function, such as equalsIgnoreCase.
@@
Recent studies (Lemieux et al., 2023; Xia et al., 2024) have demonstrated that LLMs are very effective in automated test generation. LLMs are capable of understanding program semantics and producing test inputs.
@@
Second, existing LLM-based test generation tools primarily focus on maximizing overall code coverage through automated test generation, but offer limited support for understanding which specific execution paths are exercised by each generated test. To the best of our knowledge, prior tools lack explicit test–path alignment, limiting the user’s ability to identify missing coverage and guide LLM-driven test generation toward uncovered paths.
@@ 4 String path = null;
16}
17// GPT-4o generated tests
18parseFilePath(new String[]{});
19parseFilePath(new String[]{"-v"});
20parseFilePath(new String[]{"-f", "input.txt"});
21parseFilePath(new String[]{"-v", "-f", "data.csv"});
22parseFilePath(new String[]{"-f", "log.txt", "-v"});
23parseFilePath(new String[]{"-f"});
Figure 1. This code snippet parses the argument following "-f" as a file path. While GPT-4o generates tests covering typical cases, it misses the edge case {"-f","-v"}, where "-v" is mistakenly interpreted as the file path due to the lack of validation. This path is also challenging for symbolic execution-based testing as SMT solvers like Z3 and CVC5 do not support string operations such as equalsIgnoreCase.
A Java code listing for an argument parser that returns the token following -f as the file path, plus a list of GPT-4o-generated test calls. The tests include typical cases such as {}, {-v}, and {-f, input.txt}, but do not include the input {-f, -v}.
In this paper, we present PALM, a test generation system with comprehension support that combines the path enumeration capability of symbolic execution with the test generation strength of LLMs, while sidestepping the symbolic modeling and constraint solving bottlenecks of traditional symbolic execution. PALM also features an interactive frontend that visualizes the symbolic execution tree, highlights path coverage, and allows users to inspect or refine test generation for specific paths. PALM’s test generation consists of two phases:
@@
PALM performs AST-level analysis on the subject program to systematically enumerate program paths. For each path, it automatically constructs a corresponding program variant using program transformation, which is a modified version of the original program. The variant is augmented with assertTrue and assertFalse statements that encode the branch decisions along the path. This variant provides a path-specific prompt that guides the LLM to generate the required test input. In other words, this variant directly serves as the hint for path-specific test generation.
@@
In our experiments, we evaluated PALM on 124 Java programs from HumanEval-Java (Jiang et al., 2023), which include complex control flows such as nested loops and which use external library functions. By systematically enumerating and extracting program paths, PALM achieves 35.0% and 24.2% higher path coverage than GPT-4o-mini and GPT-o3-mini, respectively, when using the same LLM backend. Its iterative test validation and refinement further improves path coverage by 14.2% over the non-iterative setting with GPT-4o-mini. In contrast, Symbolic PathFinder (a representative traditional symbolic executor) fails to derive a single accurate path constraint in 34.3% of the programs due to insufficient symbolic modeling of external API calls. We also conducted a within-subject user study with 12 participants to understand how much PALM can assist users in generating and comprehending path-aware test inputs. Participants reported higher confidence in their ability to generate sufficient tests, disambiguate redundant or missing tests, and check whether generated tests indeed exercise specific paths, compared to using a pure LLM as is.
@@
  • We conducted a user study with 12 participants, and the results show that PALM’s interactive frontend helps users better understand path coverage and identify tests that exercise specific paths.
Figure 2. PALM user interface. 1 Code editor and symbolic-execution settings. 2 Built-in example selector. 3 Start symbolic execution and test generation. 4 Symbolic execution tree (leaf nodes show coverage). 5 Select a path (click a leaf). 6 Path-specific program variant (assertions encode branch decisions). 7 Iterative test-generation history. 8 Prompt for the selected test. 9 Test editor. A Verify whether a test exercises the selected path. B Locate the corresponding path for a given test.
A screenshot of the PALM user interface annotated with numbered callouts. The interface includes a code editor and symbolic-execution settings, an example selector, a button to start symbolic execution and test generation, a symbolic execution tree with leaf nodes indicating coverage, a panel to select a path and view the corresponding path-specific program variant with embedded assertions, a history of iterative test-generation attempts, a prompt panel for the selected test, and a test editor. Additional controls allow verifying whether a test exercises the selected path and locating the path corresponding to a given test.
2. Tool Overview
@@
PALM is a test generation system for Java programs that combines symbolic execution with LLMs. For each enumerated path, PALM uses an LLM to generate a path-exercising test input. Its interface centers around a symbolic execution tree, a tree-based representation where each node corresponds to a program statement or a branch condition, and each path from the root to a leaf represents a possible execution path. The symbolic tree visualizes explored paths and highlights covered ones in green. Users can visualize the execution trace of a given test and align test inputs with path-level program execution behavior. In the following three subsections, we explain PALM’s key features using the example shown in Fig. 2.
2.1. Test generation
@@
At its core, PALM systematically enumerates program paths and invokes an LLM to generate test inputs for each enumerated path. Users begin by providing the subject program, such as a program named tutorial in the example figure ( 1 ), along with symbolic execution configurations. In this example, the method tutorial is marked as the entry point and symbolic, meaning its internal paths will be enumerated. Although the program contains no loops, users can still configure a loop bound, such as 2, to control the loop unrolling depth in programs. After clicking the Test Generation Button ( 3 ), PALM begins enumerating paths, constructing program variants, and requesting LLMs for path-specific test generation.
2.2. Path understanding with symbolic tree and path variant
@@
An interactive symbolic execution tree serves as a central component in PALM. This tree visualizes the explored execution paths of the program and distinguishes branch conditions (yellow diamond nodes) from regular statements (blue rectangular nodes). It also encodes path coverage status as green (covered), red (uncovered), and gray (unreachable). For example, the highlighted orange path ( 5 ) contains two conditional branches, x>0 and y+z>0, both of which must evaluate to true. The non-conditional assignment z=-z-5 appears between them and affects the outcome of the second condition, illustrating a data dependency along the path.
@@
The variant not only helps users interpret the logic of an execution path, but also serves as a prompt that guides the LLM to generate test inputs satisfying the conditions along that path. Unlike traditional symbolic execution tools that translate path constraints into SMT formulae, PALM leverages LLMs to handle complex language features, such as string operations and library calls, that are often inexpressible in SMT logic. This approach eliminates the need for manually modeling library functions or treating them as uninterpreted, a key advantage over conventional symbolic executors.
2.3. Validating and refining generated tests
@@
Users can inspect the divergence and hypothesize the cause. For example, the LLM might ignore the update z=-z-5 before evaluating the y+z>0 branch. To guide the LLM, users can manually edit the prompt in the Prompt Box ( 8 ) and inject guidance such as: ”The y+z>>0 branch should consider that z is updated as -z-5.” Users can also update the test to tutorial(1,6,0) and verify whether it covers the intended path ( A ), or identify the path actually taken by a test ( B ).
@@
Figure 3. PALM has two phases: (1) enumerate paths and synthesize executable path-specific variants via loop unrolling, inlining, renaming, and constant propagation/folding; (2) traverse the path tree, call an LLM to generate inputs, and iteratively validate them by execution, regenerating with feedback when a test misses the intended path.
A workflow diagram of PALM with two phases. The first phase enumerates program paths and constructs executable path-specific variants using transformations such as loop unrolling, inlining, renaming, and constant propagation and folding. The second phase traverses the resulting path tree, queries an LLM to generate test inputs for selected paths, executes the variants to validate whether the tests follow the intended path, and iteratively regenerates tests using feedback when validation fails.
3. Methodology
@@
PALM takes a Java program as input and systematically enumerates its execution paths using AST-level analysis. For each enumerated path, PALM generates a corresponding path-specific program variant, which is an executable program that represents the statements and branch decisions along the path. These variants encode each path’s semantics in a form of normal Java statements that is interpretable by LLMs, without translating to SMT constraints. To facilitate precise control over test generation, PALM organizes and visualizes paths into a symbolic execution tree.
@@
3.1. Path Extraction
@@
4 assertTrue(i < len); 2 // [1]
5 assertTrue(text.charAt(i) != // [2]
@@
4 assertTrue(i < len); 2 // [1]
5 assertFalse(text.charAt(i) != // [2]
@@
8 assertTrue(i < len); 5 // [3]
9 assertTrue(text.charAt(i) != // [4]
@@
Figure 4. Example palindrome checker with a for loop and two branches (green, yellow). (b) shows the symbolic execution tree, where each edge corresponds to a source predicate. After loop unrolling, PALM derives concrete path conditions as assertion-based encodings in (c) and (d), translating each branch to an assertTrue/assertFalse over the predicate. For example, 3 indicates the yellow predicate evaluates to true in the first iteration.
A four-part figure illustrating PALM on a palindrome checker. (a) shows a Java method is_palindrome with a for loop and a character-comparison branch. (b) shows the corresponding symbolic execution tree with edges labeled by the loop condition and the branch predicate. (c) and (d) show two path-specific program variants produced after loop unrolling, where the loop and branch predicates are encoded as assertTrue and assertFalse statements to represent different execution paths across iterations.
@@
Algorithm 1 formalizes our path-enumeration process. It recursively traverses the program’s AST and compositionally constructs the set of enumerated paths. Lines 2–5 handle If-Then-Else constructs by splitting into the two branches and inserting assertTrue or assertFalse based on the branch outcome. For example, 4 and 6 correspond to the two branches of the inner conditional check text.charAt(i) != text.charAt(len-i-1) in Fig. 4.
Lines 6–8 handle While loops by unrolling the loop body up to a user-defined loop-unrolling bound $K$. Other loop constructs such as Do-While and For loops are handled analogously. Line 10 processes compound BlockStatement structures consisting of sequential statements by enumerating the combinations of subpaths induced by each statement.
@@
0: Program programprogram, loop bound $K$ (default = 2)
[…]
1:Procedure ExtractPath(nodenode) // nodenode is an AST node
2:if nodenode is IfStatement then
3:  $P_{\text{then}}\leftarrow\texttt{PrependAll}(\texttt{assertTrue}(node.branch),\texttt{ExtractPath}(node.thenBlock))$
4:  $P_{\text{else}}\leftarrow\texttt{PrependAll}(\texttt{assertFalse}(node.branch),\texttt{ExtractPath}(node.elseBlock))$
5:  return $P_{\text{then}}\cup P_{\text{else}}$
6:else if nodenode is WhileStatement then
7:  Unroll the loop for 0 to $K$ iterations, then apply ExtractPath to each unrolled version
8:  return the union of all extracted paths
9:else if nodenode is BlockStatement then
10:  return combinations of ExtractPath($s$), for all $s\in node$
11:end if
[…]
13:End Procedure
A common practice in symbolic execution is to allow annotations specifying which functions should undergo path enumeration (referred to as symbolic functions), as well as which function serves as the entry point. This allows users to focus on path exploration on functions of interest. PALM supports this flexibility as well.
Since path extraction flattens control structures and isolates a single execution path per symbolic function, key semantic information may be lost if not preserved explicitly. To address this, we apply function inlining and variable renaming to preserve the semantics of nested calls and variable scopes. Additionally, we perform constant propagation and folding to simplify each path-specific program variant, reducing syntactic complexity and easing LLM reasoning.
@@
[…]
0:$R$: recursion bound (default = 2)
[…]
3:for recursion level $r=1$ to $R$ do
4:  $N^{\prime}\leftarrow\emptyset$
5:  for each non-terminating path $n\in N$ do
6:   Identify call-sites $c_{1},\dots,c_{k}$ in $n$ calling functions $m_{1},\dots,m_{k}$
7:   Enumerate all path tuples $(p_{1},\dots,p_{k})$ where $p_{i}\in T_{m_{i}}$
8:   $N^{\prime}\leftarrow N^{\prime}\cup$ the set of paths obtained by replacing each $c_{i}$ in $n$ with $p_{i}$
9:  end for
10:  $T\leftarrow T\cup N^{\prime}$ // Prepare for next inlining level
11:  $N\leftarrow N^{\prime}$ // Continue inlining remaining calls
12:end for
13:return $T$
@@
In the original subject program, variable scopes are managed implicitly by the nesting structure. However, after path extraction with control flow expansion, such as loop unrolling or function inlining, variable scopes are flattened. As a result, variable reuse can lead to name conflicts or ambiguity. To preserve semantic clarity and avoid confusion during test generation, PALM renames local variables by appending a numeric suffix.
@@
PALM applies intra-procedural constant propagation and constant folding. This step is performed after function inlining and variable renaming, and it only propagates numerical and boolean constants. For example, in Fig. 4, 2 contains assertTrue(i<len); where i is initialized to 0; this can be folded into assertTrue(0<len). Constant folding also enables pruning of trivially infeasible paths, such as assertTrue(x < y) where x=1 and y=0. These simplifications reduce the syntactic complexity of path variants, easing both the LLM’s symbolic reasoning and human understanding of path semantics.
3.2. Path Aware Test Generation with Validation